OpenStudy (anonymous):

Solve P = 4(3a – b) for a in terms of P and b. You must show all work.

OpenStudy (anonymous):

Distribute the 4 into the brackets first and solve by moving the terms.

OpenStudy (lopezking1):

multiply out. P = 12a - 4b Now move -4b to the left hand side because you want to isolate 12a +4b +4b So you add 4b to both sides P+4b = 12a Divide by 12 on both sides to get a P+4b/12 = a Reduce 4b/12 by dividing by 4 P+ b/___ = a Final answer

OpenStudy (anonymous):

4b / 12 = 2(2b) / (2(2)(3)) = b / 3 Cancel the two 2's in the fraction